OpenStudy (anonymous):

Graph : f(x) = (x-6)^2 + 5 Find the vertex and intercepts and axis of symmetry.

OpenStudy (callisto):

For a quadratic function f(x) = (x-h)^2 + k vertex is at (h,k) and axis of symmetry is x=h Now, h=6 and k=5. Can you get the answer for vertex and axis of symmetry? For x-int. Put f(x) =0 and solve it For y-int. Put x=0 and find the value of y. Hope it helps :)

OpenStudy (anonymous):

I'm still confused.

OpenStudy (callisto):

At which part??

OpenStudy (anonymous):

Plugging in for the x and y intercept. Do you mind starting me off?

OpenStudy (callisto):

To find x-intercept(s), put y=0, that is f(x)=0 f(x) = (x-6)^2 + 5 =0 (x-6)^2 = -5 Solve x (there should be no x-intercepts in this case, as square root of -5 is not real.) To find y-intercept, put x=0 f(0) = (0-6)^2 + 5 = (-6)^2 + 5 =?

OpenStudy (anonymous):

So what would be the x-intercept?

OpenStudy (callisto):

No x-intercepts

OpenStudy (anonymous):

Sorry I'm not the best at graphs.

OpenStudy (callisto):

See this: http://www.wolframalpha.com/input/?i=y%3D%28x-6%29%5E2+%2B+5+ => no x-intercepts

OpenStudy (anonymous):

Okay for the y-intercept?

OpenStudy (callisto):

To find y-intercept, put x=0 f(0) = (0-6)^2 + 5 = (-6)^2 + 5 = ...?

OpenStudy (anonymous):

I don't think I solved it correctly :/

OpenStudy (anonymous):

Is it 41?

OpenStudy (anonymous):

?

OpenStudy (callisto):

Yes, it is!

OpenStudy (anonymous):

So what would the line of symmetry be?

OpenStudy (anonymous):

?

OpenStudy (callisto):

axis of symmetry is x=h and h=6 So.. axis of symmetry is x= .. ?